# QRazor: Reliable and Effortless 4-bit LLM Quantization by Significant Data Razoring

### v1

@@ Abstract
Large-scale language models (LLMs) have demonstrated outstanding performance in language processing tasks, yet their deployment is often hindered by high memory demands and computational complexity. Although low-bit quantization techniques, such as 4-bit quantization, present a potential solution, they frequently lead to significant accuracy degradation or require substantial effort for such aggressive quantization approaches. To overcome these challenges, we introduce QRazor, a reliable and effortless quantization scheme designed to enable 4-bit quantization for weights, activations, and KV cache in transformer-based LLMs. The scheme involves two main stages: quantization and compression. During the quantization stage, weights, activations, and KV cache values are quantized with wider 8 or 16-bit integers as a basis to achieve nearly identical accuracy to the original full-precision LLM models, using the absolute max scaling. Subsequently, all data are compressed to 4-bit using our proposed significant data razoring (SDR) technique, which retains only the four most salient bits while discarding the others. Furthermore, we present an integer-based arithmetic unit dedicated to QRazor, enabling direct low-precision arithmetic operations without decompressing the SDR data. Despite the reduced quantization effort, QRazor achieves LLM accuracies better or comparable to state-of-the-art 4-bit methods. By also validating the hardware efficiency, our decompression-free arithmetic unit achieves 61.2% and 57.8% reduction in area and power consumption, respectively.
@@ 1 Introduction
Quantization (Dettmers et al., 2022),(Yao et al., 2022) is widely recognized as one of the most effective techniques for addressing this challenge. By aggressively reducing the precision of LLM parameters from 16-bit floating-point (FP16) to low-bit integers such as 4-bit, both memory usage and computational complexity can be significantly decreased during inference. Quantization methods are broadly divided into Post-Training Quantization (PTQ)(Banner et al., 2019) and Quantization-Aware Training (QAT)(Liu et al., 2023). This work focuses on PTQ, due to the substantial computational cost of retraining LLMs required by QAT. For low-bit integer PTQ of LLM, the key challenge is a significant accuracy drop. According to previous research (Xiao et al., 2024; Lin et al., 2024a; Yuan et al., 2023; Ashkboos et al., 2024), quantizing activations to low-bit integers is more challenging than quantizing weights due to the wider dynamic range and the presence of outliers in activations. As a result, these studies have proposed various methods to identify and mitigate activation outliers during the calibration phase of PTQ.
For instance, QLLM (Liu et al., 2024a) introduces a channel reassembly technique to minimize quantization errors in activations. However, these techniques still face significant accuracy degradation when both weights and activations are quantized to 4-bit integers. A recent study, Quarot (Ashkboos et al., 2024), successfully quantized both weights and activations to 4-bit integers while maintaining reasonable accuracy in LLMs. Notably, Quarot also demonstrated the feasibility of quantizing KV caches down to 3-bit, significantly reducing memory usage. To achieve this, Quarot utilizes orthogonal Hadamard-based rotations (Steinerberger, 2024) to suppress outliers in activations and KV caches, effectively mitigating the impact of outliers in activations. However, Hadamard-based rotation does not always guarantee effective outlier suppression despite its computation overhead. Furthermore, since the efficacy of these rotations is highly dependent on the underlying data distribution, applying Quarot to models with multiple data distributions, such as in mixture-of-expert models, requires significant tuning efforts.
This paper introduces QRazor, a PTQ method to deliver reliable 4-bit LLMs. The QRazor employs two key insights: 1) Utilizing 8-bit integers for weights and KV cache values, combined with 16-bit integers for activations, effectively captures data distributions, including outliers, and serves as the base precision scenario in this work. 2) By capturing a few salient bits from the base precision scenario, we can preserve essential characteristics of both outliers and non-outliers at reduced bit precision.
@@
This work considers the following two scenarios based on our QRazor scheme: W4A4 (4-bit weights, 4-bit activations, FP16 KV cache) and W4A4KV4 (4-bit weights, 4-bit activations, 4-bit KV cache). The accuracies of these configurations are compared against state-of-the-art (SOTA) 4-bit LLMs. For the LLaMA-1-7B and LLaMA-13B models, QRazor demonstrates superior performance, achieving more than a 10% improvement in accuracy compared to QLLM Liu et al. (2024a). When compared to Quarot, QRazor achieves significantly better results when evaluated against the rounding-to-nearest (RTN) baseline, and nearly matching results for configurations involving GPTQ for the LLaMA-2-7B and LLaMA-13B models. For other models, such as Mistral-7B and GEMMA2-2B, QRazor achieves strong accuracy results. These findings demonstrate QRazor’s ability to deliver consistent and reliable performance across diverse LLM models.
@@ 2 Related Work
3 Preliminaries: Absolute Max Scaling (ABS)
The absolute max scaling (Latotzke et al., 2022) ensures that input data fits within a target bit-width by utilizing the absolute maximum value of the data. The quantization process begins by identifying the absolute maximum value $|X_{\text{max}}|$ in the tensor, which is then used to calculate the scale factor. For example, in the case of 8-bit integer quantization, the formula to quantize a tensor $X$ under absolute max scaling and its dequantization formula are given by:
@@ 4 The QRazor Scheme
Figure 1: The overall concept of (a) typical integer-based quantization and (b) our QRazor scheme.
@@
It is important to emphasize that in the quantization stage of our QRazor scheme, activations and KV-caches utilize per-tensor and per-head scaling, respectively, while weights employ per-channel scaling. All parameters use static scaling at this stage. In contrast, Quarot, the SOTA method for quantizing all LLM parameters to 4-bit, similarly uses per-channel scaling for weights but adopts per-token scaling for activations and per-group scaling for KV caches, with a group size of 128, to enhance LLM accuracy in low-bit precision scenarios (Ashkboos et al., 2024). Notably, Quarot also relies solely on static scaling methods.
@@
Unlike weights, which are quantized offline, activations and KV caches must be quantized online using static scaling parameters. In this context, the scaling granularity of activations and KV caches plays a critical role in determining the computational complexity of LLM inference. Our coarse granularity scaling approach—per-tensor for activations and per-head for KV caches—reduces the computational complexity compared to methods employing fine granularity scaling, such as per-token for activations and per-group for KV caches in Quarot (Xiao et al., 2024). This coarse granularity effectively offsets the relative overhead introduced by compression, as discussed in the following subsections. Moreover, despite the inherent limitations of coarse granularity, our QRazor scheme achieves accuracies that are either superior or comparable to Quarot, as discussed in Section 5.1.
4.1 Quantization Stage: W8A16 or W8A16KV8 Integer Quantization
@@
In the quantization stage, LLM parameters are converted from FP16 to two base precision scenarios: W8A16 (8-bit weights, 16-bit activations) and W8A16KV8 (8-bit weights, 16-bit activations, 8-bit KV caches). These base precisions serve as primitive data types in the subsequent compression stage. Specifically, W8A16 forms the basis for W4A4, while W8A16KV8 serves as the basis for W4A4KV4, where all parameters, including KV caches, are quantized. We assume that the base precisions utilize a 2’s complement format to represent negative numbers.
As mentioned in Section 2, outliers present challenges for low-bit LLM quantization. Activations typically exhibit a more comprehensive dynamic range than weights or KV caches (Lin et al., 2024b), resulting in more extreme outliers observed within. Hence, quantizing activations effectively requires using a higher-precision format, such as 16-bit integers, to adequately capture the full range of values without significant loss of information. In contrast, we can successfully express weights and KV caches using only 8-bit integers without hardly affecting LLM accuracies.
We evaluated LLM accuracies for three scenarios: W8A8 (8-bit weights and 8-bit activations), W8A16, and W8A16KV8. These results support the hypothesis mentioned above well. Our base precision scenarios, W8A16 and W8A16KV8, achieve nearly identical accuracies to FP16-based models across various LLaMA tasks. In contrast, the quantization using W8A8 leads to significant accuracy drops, clearly demonstrating that 16-bit integers effectively characterize activation outliers, while 8-bit integers fail to do so (See Appendix A.1 for details).
  Algorithm 1 QRazor Compression of W/A/KV with Exception Handling by Group Size g.  $func$ QRazor$({X}_{i,j},w,t,g)$
Input:        ${X}_{i,j}\in\displaystyle\mathbb{R}$ — Weight, Activation, KV cache matrix,        $b_{w}$ — Quantization bit width        $b_{t}$ — Total bits to remove for compression        $g$ — Element group size Initialize:        $b_{k}=b_{w}-b_{t}$ // compressed bit width Algorithm Steps:        ${E}_{i,j}=$ Quantize(${X}_{i,j},b_{w}$)        For each element ${e}_{i,j}$ in ${E}_{i,j}$        $s_{i,j}\leftarrow$ sign bit of ${e}_{i,j}$ // extract sign bit        for $i=1,2,...,m$ do            for $j=1,2,...,n$ do                if ${e}_{i,j}<0$ do                    ${e}_{\text{si}_{i,j}}=$ remove_sign_bit(${e}_{i,j}$)                    ${e}_{\text{ci}_{i,j}}=$ concat($s_{i,j}$, 2’s_complement(${e}_{\text{si}_{i,j}}$)) // calculate 2’s complement value                else do                    ${e}_{\text{ci}_{i,j}}={e}_{i,j}$        For each element ${e}_{\text{ci}_{i,j}}$ in ${E}_{\text{ci}_{i,j}}$        ${G}_{p}=$ group_by_size_g(${e}_{\text{ci}_{i,j}},g$)        for $p=1,2,...,\displaystyle\frac{m\cdot n}{g}$ do // ($\frac{m\cdot n}{g}\in\displaystyle{\mathbb{Z}}^{+}$)            $b_{m}=$ detect_redundant_MSBs(${G}_{p},b_{t},b_{w}$) // detect redundant MSBs bit width to truncate            $b_{l}=b_{t}-b_{m}$ // calculate LSBs bit width to truncate            $F_{flagbit}=b_{l}$            ${G}_{q}=$ truncate_MSBs_and_LSBs(${e}_{\text{ci}_{i,j}},b_{m},b_{l}$) // truncate $m$ MSBs after the sign bit & truncate $l$ LSBs            if $G_{q}=2^{k}-1$ do                $G_{r}\leftarrow G_{q}^{\text{floor}}$ // truncate LSBs without carry to salient bits            else do                if MSB of truncated LSBs $=0$ do                $G_{r}\leftarrow G_{q}^{\text{floor}}$                else do                $G_{r}\leftarrow G_{q}^{\text{ceil}}$ // $r\in\{0,1,2,\dots,\frac{m\cdot n}{g}\}$, truncate LSBs with carry to salient bits         $C=\{G_{1},G_{2},\dots,G_{\frac{m\cdot n}{g}}\}$         return $C,F_{flagbit}$
@@ 4.2 Compression Stage: Significant Data Razoring
Algorithm 1 shows the pseudo-code of our QRazor scheme, where the detailed operation of the compression stage is explained. At the beginning of the compression, the base precision tensors with 2’s complement format are converted to the tensor with sign and magnitude format. This can be easily obtained by processing 2’s complement for negative values while preserving their sign bits.
Following this, our SDR technique identifies the “razoring point” for each group, which corresponds to the bit position of the leading one. The razoring point is determined by detecting the bit position of the leading one from the bitwise OR result of all data within the group (See Appendix A.3 for details). Once the razoring point is identified, a defined number of adjacent bits starting from this point, known as the salient bits, are captured. The bit width of these salient bits matches the target precision. For instance, to achieve 4-bit activations, the width of the salient bits must be four, as illustrated in Figure 1.
After capturing salient bits, we retain only the sign bit and the selected salient bits, truncating all the other bits. Then, by rounding the last bit of the remaining bits, we obtain the final compressed values. In such a scheme, during the rounding process of the least significant bits (LSBs), there is a potential risk of overflow. For example, rounding up ‘$\text{0}1111_{2}$’ where the most significant bit (MSB) represents the sign, could alter the sign bit. To prevent this, we avoid rounding the LSBs of elements where all salient bits are ‘1’. Instead, we apply flooring to the LSBs of these elements while continuing to round the LSBs of other elements within the group. The detailed procedure for handling the above exception is outlined in Algorithm 1.
@@
The size of the compression group can vary from 16 to 128. The reason we can accommodate relatively small group sizes during compression is that our SDR technique does not involve any computational quantization but simply removes zeros from the MSB portion of the integer-based data, along with a straightforward rounding process. As a result, unlike other group-based quantization approaches, there is no need for a scaling process per group using higher precision. This enables direct computation with low-bit operands, supported by our dedicated arithmetic design. The design facilitates both memory- and compute-efficient low-bit matrix multiplication, with hardware details provided in Section 4.3. It is important to note that the FP16 scaling process in QRazor is applied only per tensor for activations, per channel for weights, and per head for the KV cache.
Due to the aggressive 4-bit compression, one might be concerned that the number of zeroed elements significantly increases, degrading the reliability of our QRazor. Figure 2(a), (b) demonstrates that the first concern is effectively addressed in our QRazor. It illustrates the portion of leading ‘1’ positions before 4-bit compression (immediately after converting to the sign-and-magnitude format). The leading ‘1’ positions for activations are predominantly located between the 8th and 12th bit orders. For instance, if the leading ‘1’ position in a group is the 13th bit, parameters with an MSB below the 9th bit will be rounded and zeroed after compression, as only 4 bits are retained. Fortunately, the number of groups where the leading ‘1’ position exceeds the 12th-bit order is minimal (only 9%), indicating that outliers are infrequent and mitigating concerns about significant parameters being truncated after 4-bit compression.
In Figure2(c) we also analyzed the proportion of zeroed elements before and after 4-bit compression. For Query, Key, and Value, the increase in zeroed elements is not so considerable. However, for activations and weights, there is a notable increment. This is expected, as significant activations and weights often have small absolute values close to zero. Such truncation of small values does not lead to substantial errors. Combined with the low occurrence of outliers, as mentioned above, this explains why QRazor consistently delivers reliable performance across various LLMs, as discussed in Section 5.1.
@@
1. Eliminating Absolute Max Computation: Instead of determining the absolute maximum value within a group, QRazor detects only the leading ’1’. While the absolute maximum inherently contains the leading ’1’, multiple parameters may share the same leading ’1’ position. In DMQ, identifying the absolute maximum value which is presented in floating point data format manner, is essential for computing the scaling factor. In contrast, our compression bypasses this step entirely, significantly reducing computational complexity.
2. Lightweight Compression and Decompression: Per-group DMQ typically involves group-level quantization and dequantization operations, which rely on arithmetic computations such as multiplication and, in some cases, division. In contrast, our compression and decompression rely on bit-level truncation and shifting, which are inherently simpler and far less resource-intensive.
@@
Figure 3: The concept of decompression-free arithmetic compared with the conventional method.

### v2

@@ Abstract
Large-scale language models (LLMs) excel in language processing tasks but face deployment challenges due to high memory and computational demands. While low-bit quantization, such as 4-bit techniques, offers a potential solution, these methods often suffer from significant accuracy loss or require considerable effort for implementation such as reordering, rotation, etc. To address these challenges, we propose QRazor, a simple yet effective quantization scheme that enables 4-bit quantization of weights, activations, and KV cache in transformer-based LLMs. QRazor operates in two stages: first, quantizing data using 8 or 16-bit integers as a basis with absolute max scaling to preserve accuracy close to full-precision models, and second, compressing the quantized data to 4-bit using our significant data razoring (SDR) technique, which retains only the four most salient bits. Without any additional requirment of fine-tuning or additional training, QRazor achieves performance similar or better compared to state-of-the-art in 4-bit quantization method, surpassing Smoothquant and QLLM by over 12 points and Quarot(RTN) by more than 2.9 points in zero-shot reasoning task accuracy on the LLaMA2-7B model. Additionally, we introduce an integer-based arithmetic unit optimized for QRazor, allowing direct low-precision operations on SDR data without decompression.
@@ 1 Introduction
Quantization (Dettmers et al., 2022; Yao et al., 2022) is widely recognized as one of the most effective techniques for addressing this challenge. By aggressively reducing the precision of LLM parameters from 16-bit floating-point (FP16) to low-bit integers such as 4-bit, both memory usage and computational complexity can be significantly decreased during inference. Quantization methods are broadly divided into Post-Training Quantization (PTQ)(Banner et al., 2019) and Quantization-Aware Training (QAT)(Liu et al., 2023). This work focuses on PTQ, due to the substantial computational cost of retraining LLMs required by QAT. For low-bit integer PTQ of LLM, the key challenge is a significant accuracy drop. According to previous research (Xiao et al., 2024; Lin et al., 2024a; Yuan et al., 2023; Ashkboos et al., 2024), quantizing activations to low-bit integers is more challenging than quantizing weights due to the wider dynamic range and the presence of outliers in activations. As a result, these studies have proposed various methods to identify and mitigate activation outliers during the calibration phase of PTQ.
For instance, data redistribution techniques (Xiao et al., 2024; Liu et al., 2024a) have been widely adopted in LLMs to minimize quantization errors in activation data. However, these techniques still suffer from significant accuracy degradation when both weights and activations are quantized to 4-bit integers. Recently, a study utilizing rotation matrices (Ashkboos et al., 2024) successfully quantized both weights and activations to 4-bit integers while maintaining reasonable accuracy in LLMs. Specifically, Quarot (Ashkboos et al., 2024) utilizes orthogonal Hadamard-based rotations (Steinerberger, 2024) to suppress outliers in activations and KV caches, effectively mitigating the impact of outliers in activations. However, Hadamard-based rotation does not always guarantee effective outlier suppression despite its computation overhead. Furthermore, since the efficiency of these rotations is highly dependent on the underlying data distribution, applying Quarot to models with multiple data distributions, such as in mixture-of-expert models, requires significant tuning efforts.
This paper introduces QRazor, a PTQ method to deliver reliable 4-bit LLMs. The QRazor employs two key insights:
  • Utilizing 8-bit integers for weights and KV cache values, combined with 16-bit integers for activations, effectively captures data distributions, including outliers, and serves as the base precision scenario in this work.
  • By capturing a few salient bits from the base precision scenario, we can preserve essential characteristics of both outliers and non-outliers at reduced bit precision.
@@
This work considers the following two scenarios based on our QRazor scheme: W4A4 (4-bit weights, 4-bit activations, FP16 KV cache) and W4A4KV4 (4-bit weights, 4-bit activations, 4-bit KV cache). The accuracies of these configurations are compared against state-of-the-art (SOTA) 4-bit LLMs. For the LLaMA-1-7B and LLaMA-13B models, QRazor demonstrates superior performance, achieving more than a 10% improvement in accuracy compared to QLLM (Liu et al., 2024a). When compared to Quarot, QRazor achieves significantly better results when evaluated against the rounding-to-nearest (RTN) baseline, and nearly matching results for configurations involving GPTQ for the LLaMA-2-7B and LLaMA-13B models. For other model, such as Mistral-7B, QRazor achieves strong accuracy results. These findings demonstrate QRazor’s ability to deliver consistent and reliable performance across diverse LLM models.
@@ 2 Related Work
3 Preliminaries
Absolute Max Scaling. The absolute max scaling (Latotzke et al., 2022) ensures that input data fits within a target bit-width by utilizing the absolute maximum value of the data. The quantization process begins by identifying the absolute maximum value $|X_{\text{max}}|$ in the tensor, which is then used to calculate the scale factor. For example, in the case of 8-bit integer quantization, the formula to quantize a tensor $X$ under absolute max scaling and its dequantization formula are given by:
@@ 4 The QRazor Scheme
It is important to emphasize that in the quantization stage of our QRazor scheme, activations and KV-caches utilize per-tensor scaling, respectively, while weights employ per-channel scaling. All parameters use static scaling at this stage. In contrast, Quarot, the SOTA method for quantizing all LLM parameters to 4-bit, similarly uses per-channel scaling for weights but adopts dynamic per-token scaling for activations and per-group scaling for KV caches, with a group size of 128, to enhance LLM accuracy in low-bit precision scenarios (Ashkboos et al., 2024).
@@
Unlike weights, which are quantized offline, activations and KV caches must be quantized online using static scaling parameters. In this context, the scaling granularity of activations and KV caches plays a critical role in determining the computational complexity of LLM inference. Our coarse granularity scaling approach—per-tensor for activations and KV caches—reduces the computational complexity compared to methods employing fine granularity scaling, such as per-token for activations and per-group for KV caches in Quarot (Xiao et al., 2024). This coarse granularity effectively offsets the relative overhead introduced by compression, as discussed in the following subsections. Moreover, despite the inherent limitations of coarse granularity, our QRazor scheme achieves accuracies that are either superior or comparable to Quarot, as discussed in Section 5.2.
4.1 Quantization Stage: W8A16 or W8A16KV8
@@
In the quantization stage, LLM parameters are converted from FP16 to two base precision scenarios, depending on whether KV cache compression is applied: W8A16 (8-bit weights, 16-bit activations) and W8A16KV8 (8-bit weights, 16-bit activations, 8-bit KV caches). These base precisions serve as primitive data types in the subsequent compression stage, where they are finally converted to 4-bit integer representation.
The base precision selected for per-tensor static quantization aims to preserve outliers within the tensors without employing additional techniques such as data redistribution or rotation. It is well known that activations typically exhibit a broader dynamic range than weights or KV caches (Lin et al., 2024b), leading to more extreme outliers. Consequently, effective activation quantization requires a higher-precision format, such as 16-bit integers, to capture the full range of values while minimizing information loss.
Table 1 presents the LLM accuracy results for three different precision scenarios, applying only per-tensor static quantization. The results strongly support the need for a wider data range in activations, as the scenario quantized with 16-bit activations achieves nearly identical accuracy to its FP16 counterpart across various LLaMA tasks, whereas W8A8 leads to significant accuracy degradation. These findings highlight the necessity of 16-bit integers for capturing the outlier-prone dynamic range of activations, while 8-bit integers are sufficient for weights and KV caches. Note that our razoring scheme, applied during the compression stage, involves a precise process of detecting salient data at the bit level, directly derived from the base integer data. Therefore, selecting the appropriate base precision during the quantization stage is crucial for preserving data accuracy.
Table 1: Zero-shot accuracy of different base precision settings.
Zero-shot Accuracy $\uparrow$
Model #Bits PQ AE AC HS WG
LLaMA-2 -7B FP16 79.13 74.39 45.97 76.21 69.30
W8A8 71.89 65.99 33.11 65.11 64.33
W8A16 77.69 74.05 42.43 75.97 69.30
W8A16KV8 78.07 74.20 42.41 76.26 69.30
LLaMA-2 -13B FP16 80.23 77.82 48.76 79.39 72.30
W8A8 68.99 40.35 27.09 57.26 50.99
W8A16 78.94 76.47 45.14 78.92 71.27
W8A16KV8 78.73 76.38 44.81 79.03 71.51
@@ 4.2 Compression Stage: Significant Data Razoring
The bottom part of Figure 1 illustrates the proposed compression stage. Using sign-and-magnitude formatted integers obtained from the quantization stage, we apply our on-the-fly compression technique, termed significant data razoring (SDR). This method seamlessly integrates with the quantization stage and enables decompression-free computation, thereby minimizing the dequantization burden.
Our SDR technique identifies the “razoring point” for each group, which corresponds to the bit position of the leading one. The razoring point is determined by detecting the bit position of the leading one from the bitwise OR result of all data within the group (See Appendix A.2 for details). Once the razoring point is identified, a defined number of adjacent bits starting from this point, known as the salient bits, are captured. The bit width of these salient bits matches the target precision. For instance, to achieve 4-bit activations, the width of the salient bits must be four, as illustrated in Figure 1.
  Algorithm 1 Pseudo Code of QRazor Compression  $func$ QRazor$(X_{i,j},w,t,g)$
Input: $X_{i,j}\in\displaystyle\mathbb{R}$ — Weight, Activation, KV cache matrix, $b_{w}$ — Quantization bit width $b_{t}$ — Total bits to remove for compression $g$ — Element group size Initialize: $b_{k}=b_{w}-b_{t}$               // compressed bit width Algorithm Steps: $E_{i,j}=$ Quantize($X_{i,j},b_{w}$) For each element $e_{i,j}$ in $E_{i,j}$ $s_{i,j}\leftarrow$ sign bit of $e_{i,j}$ for $i=1,2,...,m$ do        for $j=1,2,...,n$ do           if $e_{i,j}<0$ do             $e_{\text{si}_{i,j}}=$ remove_sign_bit($e_{i,j}$)             $e_{\text{ci}_{i,j}}=$ concat($s_{i,j}$, 2’s_complement($e_{\text{si}_{i,j}}$))           else do             $e_{\text{ci}_{i,j}}=e_{i,j}$ For each element $e_{\text{ci}_{i,j}}$ in $E_{\text{ci}_{i,j}}$ $G_{p}=$ $g$($e_{\text{ci}_{i,j}},g$) for $p=1,2,...,\displaystyle\frac{m\cdot n}{g}$ do                     // ($\frac{m\cdot n}{g}\in\displaystyle Z^{+}$)        $b_{m}=$ detect_redundant_MSBs($G_{p},b_{t},b_{w}$        $b_{l}=b_{t}-b_{m}$        $F_{flagbit}=b_{l}$ $G_{q}=$ truncate_MSBs_and_LSBs($e_{\text{ci}_{i,j}},b_{m},b_{l}$)        if $G_{q}=2^{k}-1$ do        $G_{r}\leftarrow G_{q}^{\text{floor}}$               // truncate LSBs without carry        else do           if MSB of truncated LSBs $=0$ do             $G_{r}\leftarrow G_{q}^{\text{floor}}$           else do             $G_{r}\leftarrow G_{q}^{\text{ceil}}$           // $r\in\{0,1,2,\dots,\frac{m\cdot n}{g}\}$ $C=\{G_{1},G_{2},\dots,G_{\frac{m\cdot n}{g}}\}$ return $C,F_{flagbit}$
After capturing salient bits, we retain only the sign bit and the selected salient bits, truncating all the other bits. Then, by rounding the last bit of the remaining bits, we obtain the final compressed values. In such a scheme, during the rounding process of the least significant bits (LSBs), there is a potential risk of overflow. For example, rounding up “$01111_{2}$,” where the most significant bit (MSB) represents the sign, could alter the sign bit. To prevent this, we avoid rounding the LSBs of elements where all salient bits are ‘1’. Instead, we apply flooring to the LSBs of these elements while continuing to round the LSBs of other elements within the group. The detailed procedure of our QRazor scheme, including the aforementioned exception handling, is outlined in Algorithm 1.
@@
The size of the compression group is determined within the range of 16 to 128. The reason we can accommodate relatively small group sizes during compression is that our SDR technique does not involve any computational quantization but simply removes zeros from the MSB portion of the integer-based data, along with a straightforward rounding process. As a result, unlike other group-based quantization approaches, there is no need for a scaling process per group using higher precision. This enables direct computation with low-bit operands, supported by our dedicated arithmetic design. The design facilitates both memory- and compute-efficient low-bit matrix multiplication, with hardware details provided in Section 4.3. It is important to note that the FP16 scaling process in QRazor is applied only during the quantization stage, with granularities of per-tensor for activations and KV cache while per-channel for weights.
Due to the aggressive 4-bit compression, one might be concerned that the number of zeroed elements significantly increases, degrading the reliability of our QRazor. Figure 2(a), (b) demonstrates that the first concern is effectively addressed in our QRazor. It illustrates the portion of leading ‘1’ positions before 4-bit compression (immediately after converting to the sign-and-magnitude format). The leading ‘1’ positions for activations are predominantly located between the 8th and 12th bit orders. For instance, if the leading ‘1’ position in a group is the 13th bit, parameters with an MSB below the 9th bit will be rounded and zeroed after compression, as only 4 bits are retained. The empirical results show that, the number of groups where the leading ‘1’ position exceeds the 12th-bit order is minimal (only 9%), indicating that outliers are infrequent and mitigating concerns about significant parameters being truncated after 4-bit compression.
In Figure 2(c), we also analyzed the proportion of zeroed elements before and after 4-bit compression. For Query, Key, and Value, the increase in zeroed elements is not so considerable. However, for activations and weights, there is a notable increment. This is expected, as significant activations and weights often have small absolute values close to zero. Such truncation of small values does not lead to substantial errors. Combined with the low occurrence of outliers, as mentioned above, this explains why QRazor consistently delivers reliable performance across various LLMs, as discussed in Section 5.2.
@@
Eliminating Absolute Max Computation: Instead of determining the absolute maximum value within a group, QRazor detects only the leading ’1’. While the absolute maximum inherently contains the leading ’1’, multiple parameters may share the same leading ’1’ position. In DMQ, identifying the absolute maximum value which is presented in floating point data format manner, is essential for computing the scaling factor. In contrast, our compression bypasses this step entirely, significantly reducing computational complexity.
Lightweight Compression and Decompression: Per-group DMQ typically involves group-level quantization and dequantization operations, which rely on arithmetic computations such as multiplication and, in some cases, division. In contrast, our compression and decompression rely on bit-level truncation and shifting, which are inherently simpler and far less resource-intensive.
Figure 3: The concept of a commonly perceived decompression method(a) compared with the proposed decompression-free arithmetic(b).